kanchan receives a cdertian amount of money on her retirement from her employer. she give half of this money & an additional sum of Rs.10000/- to her daughter. she also gives one third of the money received & an additional sum of Rs.3000/- to her son. if the daughter gets twice as much as the son. find the amount of money kanchan received on her retirement.
Answer by HyperBrain(694)

Let x=the amount of money kanchan received on her retirement
a=amt. received by daughter
b=amt. received by son
(1)a=x/2+10000
(2)b=x/3+3000
(3)a=2b
Sltn:
Substituting (1) and (2) to (3),
x/2+10000=2(x/3+3000)
x/2+10000=2x/3+6000
6(x/2+10000)=6(2x/3+6000)
3x+60000=4x+36000
3x+60000-36000=4x+36000-36000
3x+24000=4x
x=24000
Therefore, Kanchan received Rs. 24000